What percent of 250 is 25? Include the % sign in your answer._________
Answer by Maths68(1474)

Let
25 is x% of 250

x% * 250=25
(x/100)*250=25
250x/100=25
250x=25*100
250x=2500
x=2500/250
x=10
25 is 10% of 250